1) If a consumer is currently maximizing her satisfaction, what will happen to the marginal utility of a good when its price increases? The marginal utility will____

a. Increase, because the consumer will decrease her consumption of the good

b. Decrease, because the consumer will increase her consumption of the good

c. Decrease, because the consumer will decrease her consumption of the good

d. Increase, because the consumer will increase her consumption of the good

2) the fixed cost of producing surfboards is $5,000 per month. The total variable cost of producing 15 surfboards is $36,000 per month. The average total cost of producing 15 surfboards in a month is_____.

a. $333.33

b. $2,733.33

c. $2,400.00

d. $41,000.00
Answer #1

1.

If a consumer is currently maximizing her satisfaction, what will happen to the marginal utility of a good when its price increases? The marginal utility will Increase, because the consumer will decrease her consumption of the good

option a is the correct answer

2.

TC = FC + VC = 5000 + 36000 = 41000

ATC = TC / Q = 41000 / 15 = 2733.33

option b is the correct answer